The Alpine House, Inc., is a large retailer of snow skis. The company assembled the information shown below for the quarter ended March 31:

Amount
Sales $ 1,118,000
Selling price per pair of skis $ 430
Variable selling expense per pair of skis $ 47
Variable administrative expense per pair of skis $ 16
Total fixed selling expense $ 145,000
Total fixed administrative expense $ 130,000
Beginning merchandise inventory $ 60,000
Ending merchandise inventory $ 110,000
Merchandise purchases $ 310,000

Required:

1. Prepare a traditional income statement for the quarter ended March 31.

Prepare a traditional income statement for the quarter ended March 31.

The Alpine House, Inc.
Traditional Income Statement
  
Selling and administrative expenses:
0

2. Prepare a contribution format income statement for the quarter ended March 31.

Prepare a contribution format income statement for the quarter ended March 31.

The Alpine House, Inc.
Contribution Format Income Statement
Variable expenses:
0
Fixed expenses:
0

3. What was the contribution margin per unit?
Answer #1
Units sold = Sales / Selling price per pair of skis = 1118000 / 430 = 2600
Beginning merchandise inventory 60000
(+) Merchandise purchases 310000
(-) Ending merchandise inventory 110000
Cost of goods sold 260000
Required 1 :
The Alpine House, Inc.
Traditional Income Statement
Sales 1118000
Cost of goods sold 260000
Gross margin 858000
Selling and administrative expenses:
Selling expense [ ( 2600*47 ) + 145000 ] 267200
Administrative expenses [ ( 2600*16 ) + 130000 ] 171600
438800
Net operating income 419200
Required 2 :
The Alpine House, Inc.
Contribution format Income Statement
Sales 1118000
Variable expenses :
Cost of goods sold 260000
Selling expenses ( 2600*47 ) 122200
Administrative expenses ( 2600*16 ) 41600
423800
Contribution margin 694200
Fixed expenses :
Selling expense 145000
Administrative expense 130000
275000
Net operating income 419200
Required 3 :
Contribution of each pair of skis was = Total contribution margin / Units sold = 694200 / 2600 = 267